OpenStudy (anonymous):

i need help please thanks Simplify the expression [3 + (2 + 3)2 ] • 3 102 84 42 39

OpenStudy (anonymous):

[3 + (2 + 3)2 ] • 3 u use the distributive property and get 3+4+6=13*3 because 2 times 2 is 4 and 2*3=6 and add all of them together and get thirteen and multiply by 3 to get the answer 39

OpenStudy (anonymous):

39 is the answer

OpenStudy (anonymous):

thanks